OpenStudy (anonymous):

What is the sum of the geometric sequence –4, 24, –144, … if there are 7 terms? a 308,829 b –159,964 c –411,772 d 119,973

OpenStudy (anonymous):

u know the formula ?

OpenStudy (anonymous):

Sn = a0(1-r^n)/(1-r)

OpenStudy (anonymous):

find r by dividing to successive terms . eg ->(24/-4) = -6

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

a0 = first term = -4

OpenStudy (anonymous):

what do u fail to understand ?

OpenStudy (anonymous):

I honestly just need the answer

OpenStudy (anonymous):

okay wait

OpenStudy (anonymous):

just plug in the values rebercca !

OpenStudy (paxpolaris):

\[\Large \sum_{k=1}^{7}\left( -4 \right)\left( -6 \right)^{k-1}=\huge {-4\left( 1-(-6)^7\right) \over 1-(-6)}\]

OpenStudy (anonymous):

soooo

OpenStudy (anonymous):

use ur calculator or solve manually .

OpenStudy (anonymous):

I got 223948

OpenStudy (paxpolaris):

it has to be negative...

OpenStudy (anonymous):

Just kidding I got -159964

OpenStudy (anonymous):

Thanks